OpenStudy (hannahwaller5):

Will Give Medal! The cost to produce a product is modeled by the function f(x) = 5x2 − 70x + 258 where x is the number of products produced. Complete the square to determine the minimum cost of producing this product. 5(x − 7)2 + 13; The minimum cost to produce the product is $13. 5(x − 7)2 + 13; The minimum cost to produce the product is $7. 5(x − 7)2 + 258; The minimum cost to produce the product is $7. 5(x − 7)2 + 258; The minimum cost to produce the product is $258.

OpenStudy (janu16):

5x^2-70x+258 5x^2-70x = 5(x^2-14x) = 5(x^2-14x+49 -49) = 5( x-7)^2 - 245 5x^2-70x+258 = 5(x-7)^2 -245+258 5(x-7)^2+13 The vertex is (7,13) and the parabola slopes upward because the coefficient of x^2 is positive. Therefore, (7,13) is a minimum point 5(x − 7)2 + 13; The minimum cost to produce the product is $13.5x^2-70x+258 5x^2-70x = 5(x^2-14x) = 5(x^2-14x+49 -49) = 5( x-7)^2 - 245 5x^2-70x+258 = 5(x-7)^2 -245+258 5(x-7)^2+13 The vertex is (7,13) and the parabola slopes upward because the coefficient of x^2 is positive. Therefore, (7,13) is a minimum point 5(x − 7)2 + 13; The minimum cost to produce the product is $13.

OpenStudy (hannahwaller5):

Thank you so much! That really helped! (:

OpenStudy (janu16):

yw